User Tools

Site Tools


handover:r1740

Differences

This shows you the differences between two versions of the page.

Link to this comparison view

Next revision
Previous revision
Last revision Both sides next revision
handover:r1740 [2016/05/17 17:00]
Bryn Emptage
handover:r1740 [2016/05/18 04:20]
Warren Hankey
Line 6: Line 6:
 Disk position 1799.86 GB at start. (Lucas) Disk position 1799.86 GB at start. (Lucas)
  
 +17:00:10 UT
 +2016.138.17:​00:​10 WARNING: No PCFS activity for 0.1 minutes.
 +
 +2016.138.17:​00:​10 Check VNC session to see if Field System software is running
 +
 +2016.138.17:​00:​10 Check network connection.
 +
 +Restarted field system from IP, also monica, reconfigure dbbc amd log monitor/​e-control. ​
 +
 +This lead to a new error: (e.g)
 +
 +2016.138.21:​52:​35.28?​ERROR qk -211 Tsys value for device eu overflowed or was less than zero.
 +
 +2016.138.21:​52:​39.19?​ERROR ib  -10 ibcon initialization failed, gpib devices inaccessible.
 +
 +Tried to restart GRIB from IP however this did not work.
 +
 +Jim managed to restart/​reconfigure everything; include re adjusting the attenuators which did not restore after fs reboot (is that correct Jim)?
 +
 +First '​good'​ scan:
 +138-1943 ​  755 2052-474 216 84 CW     ​19:​43:​06 ​ 19:​43:​49 ​   0:43   ​300.0 ​   *  ​
 +
 +All scans before missed. (Lucas)
 +
 +  * 23:50 UT: Manually power cycled GPIB to try and fix error ib -10 - Talk light was constantly on (possible communication jam). After power cycle - talk light is no longer constantly on. Error persists. (Bryn)
 +  * 02:06 UT: Restarted fs to try fix error ib -10. //Missed scans 139-0212 and 139-0217//. Unsuccessful - Will reboot pcfs during next gap in schedule. (Bryn)
 +  * 02:37 UT: Rebooted pcfs. (Bryn)
 +  * 02:50 UT: error ib -10 fixed, errors with dbbc communication - cycling power. (Bryn)
 +  * 03:04 UT: Reconfigured DBBC and synced time with fmset. (Bryn)
 +  * 03:10 UT: LO was not locked, restarted Rxmon. System temperatures are way off (values between -180000 and 90000) (Bryn)
 +  * 03:25 UT: Halted schedule again - trying to sort out system temperatures. (Bryn)
 +  * 04:15 UT: Schedule resumed - IF attenuation levels were not being automatically set due to a communication error (causing the errant tsys values) - First scan will be 139-0417a. **Missed/​bad scans are 139-0247bthrough 139-0415**(Bryn)
 +  * 
  
  
/home/www/auscope/opswiki/data/pages/handover/r1740.txt · Last modified: 2016/05/18 04:21 by Warren Hankey